OpenStudy (anonymous):

olve the system, using substitution. Write the solution as an ordered pair leaving no spaces, or write no solution or infinitely many. Example: (2,1) x = 2y - 2 and 3x - y = 14

OpenStudy (anonymous):

substitute the first equation into the second:\[3(2y-2)-y=14\]\[5y-6=14\]\[y=4\]Now put this y into either equation to get x:\[x=2(4)-2=6\]Solution is the point (6,4)

OpenStudy (anonymous):

but how u get 5 if you multiply by 3(2y) shouldnt it be 6y-6

OpenStudy (anonymous):

yeah. I skipped one step: 6y-6-y=14 5y-6=14...

OpenStudy (anonymous):

Ok.